OpenStudy (anonymous):

wave on a string is described by D(x,t) = 2.6cm *sin(2pi(x/8.4m) + t/(0.22s) + 1), where x is in m and t is in s. At t = 0.66s, what is the displacement of the string at x = 4.9m?

OpenStudy (anonymous):

\[D(x,t)=2.6cm*\sin(2\pi(4.9m/8.4m)+0.66s/(0.22s) +1) = \] 2.55 What am I missing here?

OpenStudy (vincent-lyon.fr):

You are probably missing an extra bracket, either here: \(D(x,t)=2.6cm*\sin(2\pi[(4.9m/8.4m)+0.66s/(0.22s)] +1) =\) or there: \(D(x,t)=2.6cm*\sin(2\pi[(4.9m/8.4m)+0.66s/(0.22s) +1]) =\)

OpenStudy (anonymous):

*facepalm* Thanks Vicent

OpenStudy (vincent-lyon.fr):

yw :)